The equation y = -x^2 + 10 models the flight of a projectile where y is the height of the projectile in meters x seconds after launch. Determine the height after:
A) 1 second
B) 3 seconds
C) 10 seconds

I know that this is probably an extremely easy problem, but I don't know where to start. I figured that you were supposed to replace a, b, and c with the x in the problem, but there are two, and I thought that would mess it up. This is for my summer packet, so there's no real rush, I just can't seem to find any online examples of this type of question. Thank you in advance!

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

The equation y = -x^2 + 10 models the flight of a projectile where y is the height of the projectile in meters x seconds after launch. Determine the height after:
A) 1 second
x = 1
h = -1 + 10 = 9 meters
--------------------
B) 3 seconds
x = 3
h = -9 + 10 = 1 meter
---------------------
C) 10 seconds
---> -90 meters. The projectile impacts before 10 seconds.
====================
y = -x^2 + 10 is not the usual equation for projectiles, tho it could be valid on some planet or asteroid.

Answer by ikleyn(52778)

.

Why you decided  (or why you think)  that the given equation models the flight of a projectile ?


            - - - > > >   It  does  not.   < < < - - -


If you don't know where to start and if you want to get familiar with the subject,  look into lessons
    - Problem on a projectile moving vertically up and down
    - Problem on an arrow shot vertically upward
    - Problem on a ball thrown vertically up from the top of a tower
    - Problem on a toy rocket launched vertically up from a tall platform

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.
Read them attentively and learn how to solve this type of problems once and for all.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Projectiles launched/thrown and moving vertically up and dawn".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.